Because of the traffic Kurt can drive home from work 20 miles faster than he can drive to work. Home from work 30 miles trip takes 1/3 less time than when he goes to work. How long does it takes Kurt to drive to work?
Answer by htmentor(1343)

Let s = the speed from home to work
Let t = the time from home to work
Since d = s*t = 30 mi, we have the following 2 equations:
30 = s*t [for the home to work trip]
30 = (s+20)*(2t/3) [work to home; 20 mph faster 1/3 less time]
Solve for s in the 1st equation and substitute in the 2nd:
(30/t + 20)(2t/3) = 20 + (40/3)t
This gives t = 3/4 hr = 45 min.